OpenStudy (anonymous):

On which of the following lines does the point (7, 1) lie? a y= 5x + 4 b y= -4x + 3 c y= -6 +8 d y= x - 10

OpenStudy (anonymous):

Test it. Let y = 1, and x =7

OpenStudy (anonymous):

1=5(7)+4? 1=-4(7)+3? 1=-6(7)+8? 1=7-10?